OpenStudy (anonymous):

show that the following progression is arithmetic and find an expression for an nth term

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+ \sqrt{x}}, \frac{ 1 }{ 1-x}, \frac{ 1 }{ 1-\sqrt{x}}\]

OpenStudy (anonymous):

@precal

OpenStudy (precal):

yuck

OpenStudy (precal):

arithmetic means that they are adding or subtracting a number to each term, so find the common difference

OpenStudy (anonymous):

yep that bit i know, tried subtracting the first term fromt the second term and subtracting the second term from the third term and the answers are different

OpenStudy (precal):